INFRARED SPECTROSCOPY OF SIZE-SELECTED PROTONATED MOLECULAR CLUSTERS: (N_{2})_{n}H^{+}, (CO)_{n}H^{+}, AND ((CH_{3})_{2}CO)_{n}H^{+}

\underline{ALLEN M. RICKS}, GARY E. DOUBERLY, and MICHAEL A. DUNCAN, Department of Chemistry, University of Georgia, Athens, Georgia 30602.

Cold, rare gas tagged, gas phase protonated molecular clusters are produced in a pulsed electric discharge supersonic expansion cluster source. The infrared spectra of the size-selected species are obtained via infrared photodissociation spectroscopy. The structure and spectroscopy (700-4000 cm^{−1}) of the monomers and dimers of protonated nitrogen, carbon monoxide and acetone will be discussed.